OpenStudy (anonymous):

what is the exponential form of 0.937

OpenStudy (rsadhvika):

\(\huge e^{\ln {.937}}\)

OpenStudy (rsadhvika):

e and ln cancel out so it is same as the original number but it is in exponent form now

OpenStudy (rsadhvika):

(.937)^1 is also an exponential form as .937 is not a perfect power of any real number i believe.. . lgbasallote sir please correct me if i am wrong

OpenStudy (anonymous):

If e^p = x, then p = ln x. Substitute: e^(ln x) = x. So the answer is just 0.937.